OpenStudy (anonymous):

Hardest math question in the world? Owen made 100 sandwiches which she sold for exactly $100. She sold caviar sandwiches for $5.00 each, the bologna sandwiches for $2.00, and the liverwurst sandwiches for 10 cents. How many of each type of sandwich did she make?

OpenStudy (anonymous):

why is this the hardest?

OpenStudy (anonymous):

is it the hardest because the girl comes from the sort of home where her parents named her owen?

OpenStudy (anonymous):

She is a sandwich

OpenStudy (anonymous):

c+b+L=100 5c+2b+0.1L=100 You have more unknowns than equations. This will have more than one answer. Too easy.

OpenStudy (anonymous):

Hehe I guess I will take that as a correct answer :D

OpenStudy (anonymous):

you may not have more than one answer as any answers that contain negatives can be thrown away (no physical meaning with respect to the question)

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Alright then. Here's a valid answer: 11 caviar sandwiches 19 bologna sandwiches 70 liverwurst sandwiches. 11+19+70=100 11*5+19*2+70*.1=55+38+7=100

OpenStudy (anonymous):

Count = zeros(101,101,101); Cost = zeros(101,101,101); Solutions = []; for i = 1:101 for j = 1:101 for k = 1:101 Count(i, j, k) = i-1+j-1+k-1; Cost(i, j, k) = 5*(i-1) +2*(j-1)+0.1*(k-1); if (Count(i,j,k) == 100) && (Cost(i,j,k)==100) newSolution = [i-1,j-1,k-1]; Solutions = [Solutions,newSolution]; end end end end give [11,19,70] for Caviar, Bologna, Liverwurst respectively

OpenStudy (anonymous):

@Mathmuse . Cool, though 8 cells and three conditions in excel with the solver did the trick for me.

OpenStudy (anonymous):

haha, dang. intially i wanted to plot the solution, but then i pulled the plug on sinking too much time into it